OpenStudy (anonymous):

When the cost of a car is multiplied by 0.08 the result is $1760 find the cost of the car, rounded to the nearest whole number. I came up with 141.00? Am I correct and can someon help me?

OpenStudy (anonymous):

Let the cost of the car is: x \[x \times (0.08) = 1760\] Divide by 0.08 both the sides to get x..

OpenStudy (cwrw238):

= 1760 / 0.08 = 176000 / 8

OpenStudy (anonymous):

so I came up with 22000. correct

OpenStudy (anonymous):

Yes dear you have done it correctly..

OpenStudy (anonymous):

Thanks waterineyes:)

OpenStudy (anonymous):

Welcome dear...

OpenStudy (anonymous):

@cwrw238 how did you come up with the answer 17600, its suppose to be divided right?

OpenStudy (anonymous):

No, he is also right.. See, I show you \[\frac{1760}{0.08} = \frac{1760 \times 100}{8} = \frac{\cancel{1760}^{220} \times 100}{\cancel{8}^1} = 220 \times 100 = 22000\]

OpenStudy (anonymous):

Ok, now i'm stuck. he came up with 17600, and mine is 22000. so what would the answer be then? sorry just want to understand if i'm doing it correctly

OpenStudy (anonymous):

See, he has written 1760/0.08 Solve this you will get the same answer..

OpenStudy (anonymous):

oh, lol. thanks a bunch

OpenStudy (anonymous):

Welcome..